OpenStudy (vipulrakta):

The power radiated by sun is 3.810 W and its radius is 710 km. The magnitude of the Poynting vector (in W/cm 2) at the surface of the sun is

OpenStudy (coacoapuffprincess):

hmmmm... One sec

OpenStudy (coacoapuffprincess):

ehhhh sorry....

OpenStudy (vipulrakta):

i need help to solve this question

OpenStudy (julyahx1):

i can help you.

OpenStudy (julyahx1):

what's the question?

OpenStudy (julyahx1):

hello?

OpenStudy (vipulrakta):

The power radiated by sun is 3.810 W and its radius is 710 km. The magnitude of the Poynting vector (in W/cm 2) at the surface of the sun is ?

OpenStudy (julyahx1):

i know the answer. But can i get a medal and fan afterwards?

OpenStudy (vipulrakta):

ok

OpenStudy (julyahx1):

The magnitude of the time-averaged Poynting vector is related to the amplitude of the electric field by S = c 1 ε E 2 . 20 y,0 So  2(1.35×103 W⋅m−2) 3 −1 Ey,0 = cε = (3.0×108 m⋅s−1)(8.85×10−12 C2 ⋅N−1 ⋅m−2) =1.01×10 V⋅m . 0 The amplitude of the magnetic field is B = E0 =1.01×103V⋅m−1 =3.4×10−6T. 0 c 3.0×108 m⋅s-1

OpenStudy (julyahx1):

medal?

OpenStudy (julyahx1):

thank you!

OpenStudy (julyahx1):

have any more questions you need help with?